A particular species of copepod, a small marine crustacean, emits a flash of light consisting of 1.4×1010 photons at a wavelength of 490 nm. The flash lasts for 2.4 s.

What is the power of the flash?
Answer #1

Energy emitted = N*h*c/lamda

= 1.4*10^10*6.625*10^-34*3*10^8/(490*10^-9)

= 5.68*10^-9 J


power of the flash = energy emiited/time

= 5.68*10^-9/2.4

= 2.37*10^-9 W